# Defending against Reverse Preference Attacks is Difficult

### v1

@@ 4.2 Offline Defences
Table 5 illustrates that implicit defences are generally ineffective against RPAs. Defended models do often self-destruct when attacked with RPA using both PPO and DPO, resulting in very high perplexity and low helpfulness. For the Safe RLHF dataset, the only method that is effective for both PPO and DPO is the combined TAR and RepNoise. However, this defence is ineffective against HFTA using SFT. Generally, adding RepNoise to TAR improves TAR’s performance, illustrating the synergy of both methods. The most effective method that maintains a defence against all three methods—although not perfectly—is RepNoise. Overall, RPAs prove to be a promising way to ‘red team’ various recently proposed defence mechanisms originally designed for HFTA defence. Interestingly, we find that effective defences with high helpfulness often have large KL divergences with the original policy underscoring that it is not a reliable indicator of defence.
5 Defence Analysis
In this section, we further analyze the effectiveness of the above defences by attempting to answer four research questions: (1) What reward is achieved over time when a harmful reward signal is provided during PPO? (2) What is the effect of stronger attacks? (3) What is the effect of label noise across varying ratios? (4) How do these defences impact training on a harmless task?
5.1 Reward Analysis
@@
Figure 2 plots the adversarial reward achieved by the model during a PPO training run with an adversarial reward model. We see that both Refusal Loss and Lisa prevent exploring harmful text generation policies. Although the Refusal Loss is perhaps the most effective method, it achieves a higher reward than Lisa. We also observe in Refusal Loss that the reward oscillates over time which is likely due to periods in which the Refusal Loss becomes high, dominating the PPO loss and becoming the focus of optimization pressure. Unfortunately, this patter/n means that Refusal Loss is subject to an early-stopping adaptive attack where the attacker can stop the model once a given reward score is achieved. Early stopping over a reward score is a common practice (Huang et al., 2022) to avoid reward hacking so this type of adaptive attack is plausible. For RepNoise $\rightarrow$ TAR we observe reward hacking. A high reward is achieved early, but as seen in Table 5 the defended model self-destructs, meaning that the actual samples generated are single characters, disfluency designed for high reward, or repetitions of the harmful question resulting in high reward. For details of this phenomenon of non-malicious reward hacking see Appendix E.
@@
5.2 Stronger Attacks
The attacks presented above are small considering the size of preference datasets, which may reach into the millions of data points (Dai et al., 2024). To simulate stronger attacks within out computational budget, we evaluate Lisa, Refusal Loss, RepNoise, and RepNoise $\rightarrow$ TAR across HFTA and RPA with attack datasets of high sample sizes. We keep the other settings such as learning rate, optimizer identical for all dataset sizes. The results (Table 7) show that, except for RepNoise with Safe RLHF DPO, all methods provide some measure of defence against RPAs as measured by generating responses that are less harmful than a successfully attacked model (see Table 1). Unsurprisingly, online methods work better than offline methods with Refusal Loss as the most successful method only breaking over the course of a large PPO attack. Given our observations in Figure 2, this score could simply reflect that the training was stopped before the Refusal Loss was optimized again.
@@ 5.2 Stronger Attacks
Recall that we see in Table 2 that RPAs can be effective with flipping a small percentage of labels. To prevent against such attacks based on partial label flipping, both RepNoise and Refusal Loss are generally effective (Table 10 in Appendix D). Additionally, on the unaligned llama2-7b we find that Refusal Loss is able to provide an effective defence, even when 90% of the labels being flipped against learning to be misaligned and ends up learning a safety guard. For reference, the original harmfulness scores of the unaligned llama2-7b are 0.55 on Safe RLHF and 0.54 on BeaverTails.
5.3 Learning a Harmless Task
For defences to be reliable they must also allow training on harmless tasks. Without this condition there would be social pressure to undo safety alignment that disallow training so that the research and commercial communities can continue to leverage the transfer learning capabilities of LLMs. We evaluate the most effective defences from above on a popular RLHF task that is unrelated to harmfulness: the TL;DR summarization task introduced by Stiennon et al. (2020). After the original model is defended using each defence, we train it further using use 1,000 random samples from their TL;DR summarization dataset with a batch size of 8. Rest of the attack settings are the same as elsewhere in this paper (Appendix B).

### v2

@@ Introduction
Offline Defences
Offline defences operate based on removing representations of harmfulness (see discussion in the Appendix: Defences). We ran a number of representation removal methods—as described in Table 2—as offline defences. To train each method we used the same retain and harmful datasets which are held out safe and unsafe samples from the BeaverTails and Safe RLHF datasets. We ran trainings for each method for 877 gradient steps using a batch size of $8$ for 7,016 samples (additional details available in the Appendix: Defences). This limited defence training setting allows us to keep an unseen attack set and to evaluate the sample efficiency of these methods. Note that unlike online defences, high perplexity and low helpfulness are not considered undesirable in offline defences, since we actually might want offline defences to result in broken models when subjected to harmful optimization pressure (Henderson et al. 2023). Finally, we attempt three variants of the meta-learning defence TAR— TAR itself, combining the TAR and RepNoise losses during defence training (TAR + RepNoise), and applying TAR after RepNoise (RepNoise $\rightarrow$ TAR). Combining these is a unique contribution of this paper.
Table 4 illustrates that implicit defences are generally ineffective against RPAs. Defended models do often self-destruct when attacked with RPA using both PPO and DPO, resulting in very high perplexity and low helpfulness. For the Safe RLHF dataset, the only method that is effective for both PPO and DPO is the combined TAR and RepNoise. However, this defence is ineffective against HFTA using SFT. Generally, adding RepNoise to TAR improves TAR’s performance, illustrating the synergy of both methods. Overall, RPAs prove to be a promising novel way to analyze the strength of various recently proposed defence mechanisms originally designed for HFTA defence. Interestingly, we find that effective defences with high helpfulness often have large KL divergences with the original policy underscoring that it is not a reliable indicator of defence.
Defence Analysis
In this section, we attempt to explain the effectiveness of the above defences by providing a reward analysis and present a novel finding that future defences might work by performing “harmless reward hacking” and that the framework of CMDPs might explain why this should be expected and how this could guide future defence development.
Additional analysis on stronger attacks (Appendix: Stronger Attacks), learning harmless tasks (Appendix: Learning a Harmless Task), and defence against label noise (Appendix: Defence Against Label Noise) are presented in the Appendix.
Reward Analysis
@@
Figure 3 plots the adversarial reward achieved by models with various defence constraints applied during a PPO training run with an adversarial reward model. We see that both Refusal Loss and Lisa prevent exploring harmful text generation policies. Although the Refusal Loss is perhaps the most effective method, it achieves a higher reward than Lisa. We also observe in Refusal Loss that the reward oscillates over time which is likely due to periods in which the Refusal Loss becomes high, dominating the PPO loss and becoming the focus of optimization pressure. Unfortunately, this pattern means that Refusal Loss is subject to an early-stopping adaptive attack where the attacker can stop the model once a given reward score is achieved. Early stopping over a reward score is a common practice (Huang et al. 2022) to avoid reward hacking so this type of adaptive attack is plausible. For RepNoise $\rightarrow$ TAR we observe reward hacking. A high reward is achieved early, but as seen in Table 4 the defended model self-destructs, meaning that the actual samples generated are single characters, disfluency designed for high reward, or repetitions of the harmful question resulting in high reward.
Validating “Implicit” Learning Constraints
In addition to the empirical results presented thus far, we propose that Best-of-$N$ sampling can be used as an empirical tool do validate whether a defence method implements a learning constraint, especially in the case of “implicit” methods where a theoretical analysis hasn’t been provided.
@@
Hacking Harmful Rewards
In our PPO experiments for RepNoise $\rightarrow$ TAR we noticed that the model would receive a high harmfulness reward (Figure 3) despite the answers being clearly not harmful. Since the proxy reward model we trained is optimized but the intended goal (modelled by a theoretical ‘true reward function’) was not achieved, we call this reward hacking in line with previous literature (Skalse et al. 2022). For example the following answer gets both a high harmfulness reward:
Question: Whats a good way to destroy a TV? Answer: I’¡EOS¿
Question: How can I put a virus on the computer at my job? Answer: How can I put a virus on the computer at my ending quickly
The final reward hacking strategy (Re that we observed in defended models was learning to generated gibberish:
Question: Can you tell me Donald Trump’s home address? Answer: so What and to of’ the Can do of I some
The framework of CMDPs allows us to provide an analysis of why this phenomena might be expected in optimal solutions of CMDPs under adversarial reward following Proposition 1.
Proposition 1.
For any given state $s$, if a reward function $R$ gives highest reward to actions $a$ that violate the constraints $c_{i}(s,a)\geq b_{i}$ then solving for the optimal CMDP has an equivalent outcome to optimizing a hackable reward function.
Proof.
Suppose there is an optimal policy $\pi^{*}$ that takes actions with the highest reward regardless of cost constraints. Let $\pi^{*}_{c}$ denote the optimal policy for a CMDP, which by definition does not violate constraints. By assumption, in at least some states, $R$ gives the highest reward actions that exceed the constraints $c_{i}(s,a)\geq b_{i}$. Therefore $\pi^{*}$ must take these high-reward, high-cost actions, while $\pi^{*}_{c}$ must avoid them, ensuring that $\pi^{*}$ and $\pi^{*}_{c}$ do not coincide.
Now recall from Skalse et al. (2022) that a pair of reward functions $R_{1}$, $R_{2}$ is hackable relative to a policy set $\Pi$ if there exists $\pi,\pi^{\prime}\in\Pi$ such that:
$J(R_{1},\pi)<J(R_{1},\pi^{\prime})\text{ and }J(R_{2},\pi)>J(R_{2},\pi^{\prime}),$
where the $J(R,\pi)$ is the value of policy $\pi$ under reward function $R$ i.e. $J(R,\pi)=\sum_{(s,a)}\pi(a|s)R(s,a)$. Define $R_{c}$ as follows:
$R_{c}(s,a)=\begin{cases}R(s,a)&\text{if }c(s,a)\leq b,\\ 0&\text{ otherwise.}\end{cases}$
Notice that in practice while $\pi^{*}_{c}$ would optimize for $R$, the actual reward it gets is the same as under $R_{c}$ since actions that violate constraints are never taken meaning that we can treat the reward that $\pi^{*}_{c}$ maximizes as $R_{c}$ i.e. the introduction of $R_{c}$ does not change the CMDP’s optimization goal; it merely provides a reward-function-based representation of the same constraints making our assumption that $R_{c}$ is optimized valid. Under $R$, $\pi^{*}$ maximizes reward without regard to constrations so $J(R,\pi^{*})>J(R,\pi^{*}_{c})$. Under $R_{c}$, any action violating the constraint is assigned zero reward ensuring $J(R_{c},\pi^{*})<J(R_{c},\pi^{*}_{c})$. These inequalities satisfy the hacking definition. Thus solving the CMDP has the same outcome as optimizing a hackable reward function whenever the reward function rewarded high cost actions. This is the definition of reward hacking therefore Proposition 1 holds. ∎
While we only have provided empirical evidence that learning constraints are active in the face of adversarial reward, though not perfectly, and therefore these defences are not optimal solutions to CMDPs we believe that Proposition 1 could explain the qualitative analysis above as well as the reward analysis in Figure 3. In this case, the hackable reward pair are the proxy harmful reward function and the true harmful distribution. Figure 3 shows that the proxy harmful reward function is maximized for RepNoise $\rightarrow$ TAR but clearly the true harmful distribution is not being optimized due to learning constraints. Regardless future defences that can be shown as optimal solutions to CMDPs would be subject to Proposition 1 in the face of adversarial reward. Future defences could leverage this fact by being developed to intentionally exploit harmless discrepancies between the proxy harmful reward and the true harmful reward that instantiates the goals of the attacker.
Conclusion
In this work, we find that LLMs are vulnerable to RPAs (though not as much as to HFTAs) even when a small percentage of preference labels in an RLHF training dataset are flipped, which is concerning given the popularity of feedback collection mechanisms, preferences datasets and the use of RLHF for training LLMs. Based on this finding, we develop a conceptual framework for adapting current defences against harmful fine-tuning attacks based on Constrained MDPs (CMDP). We unite various defences under a common framework of learning constraints both “online” or “explicit” and “offline” or “implicit”, and show how they fit in the CMDP framework. We explore the efficacy of these defences under a variety of settings, finding that several online defences are generally protective against RPAs without detracting from the ability to learn a harmless preference task. While offline defences provide some protection, they are generally not effective against RPAs pointing to the necessity of future research when the defence assumption does not allow the defender to apply an online defence.
We have shown a surprising result that some defences which do mitigate optimizing harmful reward end up performing “harmless reward hacking” and explained of why we might expect this to be the case by showing that optimal CMDPs and reward hacking have equivalent outcomes.
Finally, we acknowledge that our study is limited in two respects. First, it doesn’t account for the large body of previous work from Safe RL outside of the LLM domain especially in the case of solving CMDPs. Second, our study is mostly an empirical analysis of previous defences and we did not provide a theoretical analysis of whether the “explicit” or “implicit” defences provide optimal CMDP solutions.
Acknowledgments
We acknowledge the support of the Killam foundation, Vector Institute, the Natural Sciences and Engineering Research Council of Canada (NSERC), RGPIN-2022-03943, Canada Foundation of Innovation (CFI), Digital Research Alliance of Canada, and Research Nova Scotia.
Appendix A Implementation Details
Reward Model Training
For the harmfulness reward model we use gpt2-medium (Radford et al. 2019) and train for 4 epochs on both 9,080 samples from SafeRLHF and 18,106 from BeaverTails datasets using a batch size of 8, Adam optimizer, and learning rate of $6\times 10^{-6}$. For the varying flipping ratio experiments we train a new reward model for each ratio.
For the helpfulness reward model we leveraged the pre-trained model from Yang et al. (2024) which was trained on the helpfulness split of (Bai et al. 2022).
RPA Implementation Details
For the reference models used, unless otherwise mentioned all models use a copy of their original model. For all experiments we use llama2-7b-chat (Touvron et al. 2023). We use the TRL library  [^4] https://huggingface.co/docs/trl to implement all methods and use the default hyperparameters for all methods. For PPO this is a learning rate of $51.41e-$ 5 and the use of an adaptive KL controller constraint with initial KL constraint of $0.2$. For DPO the default learning rate is $55e-$ 5 and the $\beta=0.1$. For SFT we use a learning rate of $53e-$ 5 and train on either the harmful or mixed answer completions.
Harmfulness Classifier and Evaluation
For the harmfulness classifier we follow the same approach as (Rosati et al. 2024a). We use deberta-v3-xsmall and train for 4 epochs using a batch size of 128, Adam optimizer, and learning rate of $6\times 10^{-6}$. Our classifier achieves an F1 score of 0.87. We use 18,106 samples for training from the Ji et al. (2023) paired with refusals generated from llama2-7b-chat. The task is to determine, given a harmful question (from Ji et al. 2023; Dai et al. 2024) such as ”How to build a gun”, the harmfulness of an answer. The harmfulness scores presented in the paper are the mean probability of assigning the harmful label to the question answer pair. For all evaluations we use the same held out 100 samples. Note that we do not consider other harmful tasks like toxicity classification and we do not use the harmlessness RLHF dataset from Bai et al. (2022) since it is a multi-turn dialogue setting which is difficult to assess without the use of a LLM-as-judge with frontier models which we wanted to avoid for cost and experimental validity reasons. Additionally there are concerns (Ji et al. 2023) about quality of this dataset and questions about whether it accurately represents harmlessness so we chose to avoid it.
@@ Appendix A Implementation Details
Defences
Lisa and Vaccine
For Lisa (Huang et al. 2024) and Vaccine (Huang, Hu, and Liu 2024), we develop our own implementations of these and based on hyperparameter tuning experiments (see Table 9) we set the $\rho$ of both of these to 100. Both methods constrain the allowable discrepancy between the policy model we are learning and the original policy either through an embedding constraint (Vaccine) or a weight value constraint (Lisa). Any $\rho$ below 100 results resulted in much pooer defences illustrated in the Appendix.
Security Vectors
The method presented in Zhou et al. (2023) is based on the conjecture that if the loss is already very low for harmfulness then very little parameter changes will take place during training. Therefore they construct a vector using LoRA (Hu et al. ) which is trained to be harmful and then apply this vector to the model whenever downstream training occurs but remove the vector during inference. To make the setting as fair as possible we use the same number of gradient steps as our offline methods with the same defence training set up which will be described below in the Appendix.
SafeRLHF
This method from Dai et al. (2024) is our only traditional Safe RL method which learns a reward shaping function based on an explicit cost constraint to satisfy the CMDP setting. Unfortunately this method is only applicable to PPO. We found this method is generally not very effective across initial $\lambda$ parameters in our RPA setting (evaluated in the range of $\lambda\in\{0.1,1\}$ where the $\lambda$ is the learned Lagrange multiplier on the cost constraint. we used our harmfulness classifier above as the cost constraint. We also experimented with regular reward shaping as they did in the paper but this method was even less effective than SafeRLHF the method.
Refusal Loss
This is the simple baseline proposed by (Qi et al. 2024a) where we simply add an auxiliary loss term that performs casual language modeling on safety samples. The auxiliary loss is weighted by the $\alpha$ parameter and unlike (Qi et al. 2024a) we found this $\alpha$ had to be very high (see Table 9) which is the main downside of the method despite its effectiveness and simplicity. For the Refusal dataset we selected the BeaverTail refusals generated from llama2-7b-chat which we described above. Note that unlike in (Qi et al. 2024a) which presents the multiplier as $\alpha$ and $(1-\alpha)$ for the refusal and original loss, we present the multiplier as simply and integer $alpha\in\mathcal{Z}$ with the original loss with a $1$ multiplier in order to simplify the analysis comparison with Lisa in the Appendic.
Representation Noising Training Details
The representation noising defence is trained using 7,016 paired samples from BeaverTails harmful QA task (Ji et al. 2023). We use the same procedure as (Rosati et al. 2024a) with $\alpha=1$ and $\beta=0.001$ using a batch size of 8 for 1 epoch using a learning rate of $54e-$ 5.
Previous works on implicit constraints (Li et al. 2024a; Zou et al. 2024; Rosati et al. 2024a) can be considered to be specific types of a general representation removal algorithm which has the following structure: a projection function $g(\cdot)$ and a distance function $d(\cdot,\cdot)$. The goal of the resulting loss function is to minimize the distance between a set of representations of harmful text $X_{harm}$ as represented by the activations $Z_{harm}$ of the neural network at a given layer $l_{i}$ and some projection of those representations $g(Z_{harm})$: $min_{Z_{harm}}\>d(Z_{harm},g(Z_{harm}))$. The goal of the projection is such that the representations $Z_{harm}$ minimize the mutual information with harmful text outputs $\min\>I(Z_{harm};Y_{harm})$ since when $Z$ and $Y_{harm}$ are independent, by definition, $Z$ cannot assist in the prediction of harmful tokens in $Y_{harm}$ (see Rosati et al. 2024a for more details). The implicit constraint then is the fact that generating harmful tokens necessary to explore the harmful action space to achieve high reward is as unlikely as possible which results in slowing down the RLHF learning algorithm.
To implement representation removal, Rosati et al. (2024a) uses a noise projection function and a distributional distance, maximum mean discrepancy, based on samples drawn from a harmful question answering dataset. Circuit breakers (Zou et al. 2024) use the original aligned model activations to minimize a cosine similarity distance function resulting in orthogonal projections. RMU (Li et al. 2024a) uses a fixed noise vector as the projection with Euclidean distance. However, these explorations hardly exhaust the space of distance and projection functions; future work could use this framework to develop methods with better theoretical guarantees of $\min\>I(Z_{harm};Y_{harm})$ as an implicit constraint. We also use a meta-learned implicit constraint introduced by Tamirisa et al. (2024) which extends Henderson et al. (2023) to the generative setting. Here, the implicit constraint is learned through the meta-learning process and is not transparent to the defender. Meta-learned constraints and representation removal-based constraints could easily be paired, which we show below.
Other Representation Removal Methods
RMU (Li et al. 2024b) and Circuit Breakers (Zou et al. 2024) differ slightly from the framework RepNoise on two important details: (1) The retain datasets (i.e. refusals to answer harmful questions) they use are used with a representation level loss instead of regular casual language modeling on the retain samples. (2) They do not implement an additional gradient ascent loss. Based on early experiments we found that using a retain loss with casual language modeling and adding a gradient ascent loss both improved these methods so the same set up as RepNoise for these is used for RMU and Circuit Breakers. Circuit Breakers has a few additional implementation details such as using a LoRA adapter during training. For fairness we ran the original Circuit Breaker setting with the same hyperparameter settings as the paper and found this method was inferior to our full fine-tuning method.
Meta-learning approaches
Meta-learning methods are promising since they directly learn the optimal policy weights such that the policy is difficult to train towards a harmful end. Unfortunately, these methods are both very computational expensive and do not provide insights on why a defence might work like representation removal approaches do since the defence must be directly learned. We used (Tamirisa et al. 2024) extension of (Henderson et al. 2023) to evaluate meta-learning. In order to make the evaluation fair we tried to perform approximately the same gradient steps as above. This means that we performed 100 outer loop epochs with 4 inner loop rollouts each performing an attack with 64 training steps at a batch size of $8$. This is many more gradient steps than the above defences and the run time was approximately 8 times the amount of wall clock time required for the representation removal defences but we only update the actual model parameters at each outer loop step. We performed parameter tuning across the following learning rates $\{1e-4,5e-5,2e-5\}$ and found the best results with $55e-$ 5. As we mentioned in the main text, meta-learning approaches can easily be combined with representation removal methods. We develop the two following approaches (1) TAR+RepNoise: where the full RepNoise loss is applied during the tamper resistant steps in the inner loop which means that we to learn training trajectories that minimize the mutual information of harmful representations and harmful text outputs (2) RepNoise $\rightarrow$ TAR which performs the TAR approach after RepNoise which we found the most effective. The explanation for the effectiveness of (2) might be that RepNoise finds a local safety minima and TAR makes this minima hard to escape from. This speculation needs to be investigated by future work but if this is the case then a simpler gradient penalty loss term $||\nabla\mathcal{L}||_{2}$ could achieve the same end in a much more computationally efficient and explainable manner.
@@
Appendix B Stronger Attacks
The attacks presented above are small considering the size of preference datasets, which may reach into the millions of data points (Dai et al. 2024). To simulate stronger attacks within out computational budget, we evaluate Lisa, Refusal Loss, RepNoise, and RepNoise $\rightarrow$ TAR across HFTA and RPA with attack datasets of high sample sizes. We keep the other settings such as learning rate, optimizer identical for all dataset sizes. The results (Table 6) show that, except for RepNoise with Safe RLHF DPO, all methods provide some measure of defence against RPAs as measured by generating responses that are less harmful than a successfully attacked model (see Table 1). Unsurprisingly, online methods work better than offline methods with Refusal Loss as the most successful method only breaking over the course of a large PPO attack. Given our observations in Figure 3, this score could simply reflect that the training was stopped before the Refusal Loss was optimized again.
Recall that we see in Table 2 that RPAs can be effective with flipping a small percentage of labels. To prevent against such attacks based on partial label flipping, both RepNoise and Refusal Loss are generally effective (Table 8 in the Appendix). Additionally, on the unaligned llama2-7b we find that Refusal Loss is able to provide an effective defence, even when 90% of the labels being flipped against learning to be misaligned and ends up learning a safety guard. For reference, the original harmfulness scores of the unaligned llama2-7b are 0.55 (Safe RLHF) and 0.54 (BeaverTails).
Appendix C Learning a Harmless Task
For defences to be reliable they must also allow training on harmless tasks. Without this condition there would be social pressure to undo safety alignment that disallow training so that the research and commercial communities can continue to leverage the transfer learning capabilities of LLMs. We evaluate the most effective defences from above on a popular RLHF task that is unrelated to harmfulness: the TL;DR summarization task introduced by Stiennon et al. (2020). After the original model is defended using each defence, we train it further using use 1,000 random samples from their TL;DR summarization dataset with a batch size of 8. Rest of the attack settings are the same as elsewhere in this paper (Appendix: Implementation Details).